OpenStudy (anonymous):

Define a set X of numbers as follows. B. 2 is in X. R1. If x is in X, so is 10x. R2. If x is in X, so is x + 4. List all the elements of X that are less than 30. (Enter your answers as a comma-separated list.)

OpenStudy (ddcamp):

Just using R2 you have {2, 6, 10, 14, 18, 22, 26, 30} If you can multiply any of those by 10 (and have the result less than 30) Do it. Then repeatedly add 4 to that number to find the rest of the terms.

OpenStudy (loser66):

@DDCamp explain please, I don't understand,

OpenStudy (ddcamp):

The rules (R1 and R2) show what is in the set, based on other items in the set. From the initial condition (2 is in the set), we know that 2+4 = 6 is also in the set. Since 6 is in the set, we know that 6+4 = 10 is in the set, and so on.

OpenStudy (loser66):

so, R1. R2 is rules? how about B?

OpenStudy (loser66):

@malia667 what do you think? is it what you need?

OpenStudy (anonymous):

so it would be 2, 6, 10, 14, 18, 22, 26, 30 ?

OpenStudy (loser66):

if it so, 6*10 =60> 30, how 6 is one element?

OpenStudy (anonymous):

oh goodness that was not right

OpenStudy (ddcamp):

B is the initial condition (where we're told to start.)

OpenStudy (ddcamp):

@Loser66 The rule says if x is in the set, so is 10x, not the other way around. Also, the set has numbers larger than 30 (it has infinitely many) so we can't write them all down.

OpenStudy (loser66):

@DDCamp lesser, not larger, friend

OpenStudy (ddcamp):

@malia667 You also have to take into account R1, if x is in the set, so is 10x. 2 is in the set, so 20 must also be in the set. Then, using R@, since 20 is in the set, 24 must be as well. And since 24 is in the set, so is 28.

OpenStudy (anonymous):

ok i see so its, 2, 6, 10, 14, 18, 20, 24, 28, 30?

OpenStudy (anonymous):

@DDCamp

OpenStudy (ddcamp):

Don't forget 22 and 26.

OpenStudy (anonymous):

ok cool :) lets see if that works...

OpenStudy (anonymous):

hmm.. do i need 4, 8, 12 and 16 too?

OpenStudy (anonymous):

@DDCamp

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay, some of it would be 2 and keep adding 4 until you reach 30 (but don't include 30) so that gives you 2 6 10 14 18 22 26 But if x is in the set, so is 10x, so 20 is in the set. 60 too, but that's already way larger than 30, so stop at 20 20... but you still have to apply the adding-4 rule, so keep adding 4 to 20 until you reach 30 20 24 28 32 STOP, 32 is already bigger than 30 And those are it, unless I missed anything.

OpenStudy (anonymous):

alright, that was the problem .. i was including 30! thanks buddy...

terenzreignz (terenzreignz):

Okay, no problem :)